OpenStudy (anonymous):

(cos5-√3.sin5)/sin25=?

OpenStudy (anonymous):

\[\cos 5-\sqrt{3} \ \sin 5=2(\frac{1}{2} \cos 5 - \frac{\sqrt{3}}{2} \sin5)=2(\sin 30 \cos 5- \cos 30 \sin 5)\]

OpenStudy (anonymous):

now what will u get?

OpenStudy (anonymous):

Not more

OpenStudy (anonymous):

ı think